# Operator Learning for Continuous Spatial-Temporal Model with Gradient-Based and Derivative-Free Optimization Methods

### v2

@@ 2.2 Neural Operator
To construct the data-driven dynamical operator $\tilde{\mathcal{G}}$ in the continuous setting, i.e., viewing the system state $\bm{u}$ as a continuous spatial-temporal field instead of a discrete finite-dimensional vector, we rely on the recent developments of neural operators li2021fourier ; lu2021learning . More specifically, we choose to use Fourier neural operator (FNO) li2021fourier as a tool to construct the data-driven dynamical operator $\tilde{\mathcal{G}}$ and would like to point out that other neural operator architectures (e.g., DeepONet  lu2021learning ) can also be used in the framework of this paper. In general, neural operators aim to approximate a non-linear mapping between infinite-dimensional spaces:
$\mathcal{G}:\mathcal{A}\mapsto\mathcal{B},$ (2.5)
with a neural network $\tilde{\mathcal{G}}(\cdot;\theta)$ parameterized by $\theta$. $\mathcal{A}=\mathcal{A}(D;\mathbb{R}^{d_{\mathrm{a}}})$ and $\mathcal{B}=\mathcal{B}(D;\mathbb{R}^{d_{\mathrm{b}}})$ be separable Banach space of function defined on $D\subset\mathbb{R}^{d}$ be a bounded open set. The key advantages of continuous mapping in Eq. (2.5) are: (i) the performance of the trained mapping is resolution-invariant, and (ii) the flexibility of using data with different discretizations.
To approximate the continuous mapping in Eq. (2.5) with a neural operator, the standard FNO framework assumes that we have observed data pairs $\{\mathrm{a}_{j}|_{D_{j}},\mathrm{b}_{j}|_{D_{j}}\}$, where $\mathrm{a}_{j}$ are i.i.d. samples from its probability measure $\mathbb{P}$, and $D_{j}=\{x_{1},...,x_{n}\}\subset D_{\bm{x}}$ is a $n$-point discretization of the domain $D_{\bm{x}}$. It should be noted that $\mathrm{a}_{j}|_{D_{j}}\in\mathbb{R}^{n\times d_{\mathrm{a}}}$ denotes the evaluation of $\mathrm{a}_{j}$ on the set of discrete points $D_{j}$. In this work, we focus on the use of FNO to construct the data-driven operator $\tilde{\mathcal{G}}$, which first linearly transforms input function evaluation $\mathrm{a}(x)$ to lift the dimension from $\mathbb{R}^{d_{\mathrm{a}}}$ to $\mathbb{R}^{d_{v}}$: $v_{0}(x)=P(\mathrm{a}(x))$. Then $v_{0}(x)$ will be iteratively transformed by Fourier layers with the output dimension staying the same as $d_{v}$:
$v_{n+1}(x)=\sigma(Wv_{n}(x)+(Kv_{n})(x)),\quad n=0,1\ldots N-1,$ (2.6)
@@
where $(Kv_{n})(x)=\mathcal{F}^{-1}(R\cdot(\mathcal{F}v_{n}))(x)$ is the Fourier integral operator, $\mathcal{F}$ is Fourier transform and $\mathcal{F}^{-1}$ is its inverse, $R$ is the complex-valued parameters in the neural network, and $W$ is a linear transformation. At last, $v_{N}(x)$ will be linearly transformed again to ensure the dimension of the final output same as the solution function in the space $\mathcal{B}$, i.e., $\tilde{\mathrm{b}}(x)=\tilde{\mathcal{G}}(\mathrm{a};\theta)(x)=Qv_{N}(x)\in\mathbb{R}^{d_{\mathrm{b}}}$.
The parameters of Fourier neural operator $\tilde{\mathcal{G}}$ can be optimized by:
$\min_{\theta}\mathbb{E}_{\mathrm{a}\sim\mathbb{P}}[L(\mathcal{G}(\mathrm{a};\theta),\tilde{\mathcal{G}}(\mathrm{a};\theta))],$ (2.7)
where $L:\mathcal{B}\times\mathcal{B}\mapsto\mathbb{R}$ is a loss functional and is often estimated based on the set of discrete points $D_{j}$.
@@
In this work, we aim to construct the dynamical operator $\tilde{\mathcal{G}}$ in Eq. (2.1) via FNO. However, the data of $\partial\bm{u}(\bm{x},t)/\partial t$ may not be available in many applications, e.g., when the temporal resolution in the data of $\bm{u}(\bm{x},t)$ is not fine enough to obtain an informative estimation of its time derivative. Therefore, we would not be able to directly train the data-driven dynamical operator $\tilde{\mathcal{G}}$ by solving the optimization problem in Eq. (2.7). To address this issue, we explore the training of a neural dynamical operator via the framework of neural ODE chen2018neural .
@@ 2.3 Neural ODE
The neural ODE framework chen2018neural focuses on a general form of dynamical system:
$\frac{\mathop{}\!\mathrm{d}\bm{z}(t)}{\mathop{}\!\mathrm{d}t}=\bm{f}(\bm{z}(t),t),$ (2.8)
where $\bm{z}(t)\in\mathbb{R}^{d_{\bm{z}}}$, $t\in[0,T]\subset\mathbb{R}$, and $\bm{f}:\mathbb{R}^{d_{\bm{z}}}\times[0,T]\mapsto\mathbb{R}^{d_{\bm{z}}}$. The goal of neural ODE is to train a surrogate model $\tilde{\bm{f}}$ to approximate $\bm{f}$ based on a time series data of $\bm{z}(t)$. The surrogate model is often constructed by a neural network $\tilde{\bm{f}}(\bm{z},t;\theta)$ parameterized with trainable parameters $\theta$.
In our work, $\bm{z}(t)$ corresponds to $\bm{u}_{t}|_{D_{j}}$, which denotes the evaluation of $\bm{u}(\cdot,t)$ in Eq. (2.1) on a set of discrete points $D_{j}$. With a given $\tilde{\bm{f}}$ and the true system state at $t_{i}$, the simulated system state at a future time $t_{i+1}$ can be written as $\tilde{\bm{z}}(t_{i+1})=\bm{z}(t_{i})+\int_{t_{i}}^{t_{i+1}}\tilde{\bm{f}}(\tilde{\bm{z}}(t),t;\theta)\mathop{}\!\mathrm{d}t$, which can be obtained by an ODE solver in real applications.
Given observed states $\{\bm{z}({t_{i}})\}_{n=0}^{N}$, the trainable parameters $\theta$ in $\tilde{\bm{f}}$ can be optimized via:
$\min_{\theta}L:=\sum_{i=0}^{N}\|\bm{z}(t_{i}),~{}\tilde{\bm{z}}(t_{i})\|^{2},$ (2.9)
@@
noting that $\tilde{\bm{z}}(t_{i})$ depends on $\theta$ for all $i>0$. The loss function $L$ is often chosen as the standard vector $\ell^{2}$-norm. The neural ODE framework chen2018neural discussed two methods to calculate $\frac{\mathop{}\!\mathrm{d}L}{\mathop{}\!\mathrm{d}\theta}$, including (i) backpropogation and (ii) adjoint sensitivity method, with which we can solve the optimization problem in Eq. (2.9) via gradient descent.
@@
Backpropogation is a classical gradient descent optimization method for training a neural network. One drawback of the backpropogation for neural ODE training is the memory cost to store the results of its forward pass, which is linearly proportional to the number of $\tilde{\bm{f}}$ evaluations. Although various types of memory management techniques margossian2019review have been developed for automatic differentiation, the memory cost for matching a long time series data with a large scale model of $\tilde{\bm{f}}$ can still be inefficient or even infeasible.
@@
On the other hand, the adjoint sensitivity method calculates the gradient $\frac{\mathop{}\!\mathrm{d}L}{\mathop{}\!\mathrm{d}\theta}$ by a forward and a backward ODE integration. The forward integration solves for the state $\tilde{\bm{z}}$ at $t_{1},t_{2},...,t_{N}$, with the initial condition $\tilde{\bm{z}}(t_{0})=\bm{z}(t_{0})$:
$\frac{\mathop{}\!\mathrm{d}\tilde{\bm{z}}(t)}{\mathop{}\!\mathrm{d}t}=\tilde{\bm{f}}(\tilde{\bm{z}}(t),t;\theta).$ (2.10)
By introducing an adjoint state $\bm{a}(t_{i})=\frac{\partial L}{\partial\tilde{\bm{z}}(t_{i})}$, the backward integration solves for an augmented states $[\tilde{\bm{z}}(t),\bm{a}(t),\frac{\mathop{}\!\mathrm{d}L}{\mathop{}\!\mathrm{d}\theta}]^{\top}$ from $t_{i}$ to $t_{i-1}$, for $i=N,N-1,...,1$, with the initial state $\tilde{\bm{z}}(t_{N})$ evaluated from the forward integration above, $\bm{a}(t_{N})=\frac{\partial L}{\partial\tilde{\bm{z}}(t_{N})}$, and $\frac{\mathop{}\!\mathrm{d}L}{\mathop{}\!\mathrm{d}\theta}|_{t_{N}}=\bm{0}$:
$\frac{\mathop{}\!\mathrm{d}\tilde{\bm{z}}(t)}{\mathop{}\!\mathrm{d}t}=\tilde{\bm{f}}(\tilde{\bm{z}}(t);\theta),$ (2.11)
$\frac{\mathop{}\!\mathrm{d}\bm{a}(t)}{\mathop{}\!\mathrm{d}t}=-\bm{a}(t)^{\top}\frac{\partial\bm{\tilde{f}}(\tilde{\bm{z}}(t),t;\theta)}{\partial\tilde{\bm{z}}(t)},$
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\frac{\mathop{}\!\mathrm{d}L}{\mathop{}\!\mathrm{d}\theta}=-\bm{a}(t)^{\top}\frac{\partial\bm{\tilde{f}}(\tilde{\bm{z}}(t),t;\theta)}{\partial\theta}.$
@@
It should be noted that the solved adjoint state $\bm{a}(t_{i})$ needs to be adjusted at $i=N-1,N-2,...,0$ during the backward integration, by adding a term of $\frac{\partial L}{\partial\tilde{\bm{z}}(t_{i})}$ to account for the fact that the loss function $L$ explicitly depends on the system state $\tilde{\bm{z}}(t_{i})$. During the whole backward integration, the vector-Jacobian products $\bm{a}(t)^{\top}\partial\bm{\tilde{f}}/\partial\tilde{\bm{z}}$ and $\bm{a}(t)^{\top}\partial\bm{\tilde{f}}/\partial\theta$ can be evaluated on the fly without storing them in the memory. Therefore, the adjoint sensitivity method has a constant memory usage $O(1)$ with respect to the integration time steps, i.e., the number of $\tilde{\bm{f}}$ evaluations.
@@
For a short-term integration, the backpropogation often has no memory cost issue and is more computationally faster than adjoint method, considering that the Jacobian is stored and does not need to be evaluated on the fly in the backward pass. However, the memory cost issue would prevent the use of backpropogation if the data involves long-term integration, e.g., time-averaged statistics. In addition, gradient-based optimization can potentially encounter numerical issues (e.g., gradient blow-up) for the long-term information of chaotic systems. Therefore, backpropogation is used in this paper for short-term trajectory matching, while the long-term statistics of a chaotic dynamical system is incorporated by a derivative-free Kalman method, instead of adjoint sensitivity method. More details about the derivative-free Kalman method can be found in section  2.4.
@@ 2.4 Ensemble Kalman Inversion
Originated from ensemble Kalman filter evensen1994sequential ; evensen2003ensemble ; evensen2009data , ensemble Kalman inversion (EKI) iglesias2013ensemble ; kovachki2019ensemble has been developed as an optimization method to solve inverse problems. More specifically, the goal of EKI is to estimate the unknown parameters from noisy data in the general form of an inverse problem:
$y=G(\theta)+\eta$ (2.12)
where $y\in\mathbb{R}^{d}$ is a vector of observation data, $G$ denotes a forward map, $\theta\in\mathbb{R}^{p}$ represents the unknown parameters, $\eta\sim\mathcal{N}(0,\Sigma_{\eta})$ is often chosen as Gaussian random noises with a covariance matrix $\Sigma_{\eta}\in\mathbb{R}^{d\times d}$.
The goal of EKI is to identify the optimal parameters $\theta$ that minimize the loss function:
$\Phi_{\theta}=||\Sigma_{\eta}^{-\frac{1}{2}}(y-G(\theta))||^{2},$ (2.13)
which corresponds to the negative log-likelihood under the Gaussian distribution assumption. The formula that iteratively updates the ensemble parameters $\{\theta^{(j)}\}_{j=1}^{J}$ is:
$\begin{gathered}\theta_{n+1}^{(j)}=\theta_{n}^{(j)}+\Sigma_{n}^{\theta g}(\Sigma_{n}^{gg}+\Sigma_{\eta})^{-1}(y^{(j)}-g_{n}^{(j)})\end{gathered}$ (2.14)
where the index $n$ denotes the $n$-th EKI step and $y^{(j)}$ corresponds to the perturbed observation data $y$ via sampling the noises $\eta$. With the ensemble of parameters $\{\theta_{n}^{(j)}\}_{j=1}^{J}$ at the $n$-th EKI step, the terms $g_{n}^{(j)}$, $\Sigma_{n}^{\theta g}$, and $\Sigma_{n}^{gg}$ in Eq. (2.14) are calculated as:
$\begin{gathered}\bar{\theta}_{n}=\frac{1}{J}\sum_{j=1}^{J}\theta_{n}^{(j)},~{}~{}g^{(j)}_{n}=G(\theta_{n}^{(j)}),~{}~{}\bar{g}_{n}=\frac{1}{J}\sum_{j=1}^{J}g_{n}^{(j)},\\ \Sigma_{n}^{\theta g}=\frac{1}{J-1}\sum_{j=1}^{J}(\theta_{n}^{(j)}-\bar{\theta}_{n})(g_{n}^{(j)}-\bar{g}_{n})^{T},\\ \Sigma_{n}^{gg}=\frac{1}{J-1}\sum_{j=1}^{J}(g_{n}^{(j)}-\bar{g}_{n})(g_{n}^{(j)}-\bar{g}_{n})^{T}.\end{gathered}$ (2.15)
In this paper, we utilize EKI to solve the optimization problem Eq. (2.4). The forward map $G$ is a composition of the data-driven dynamical operator $\tilde{\mathcal{G}}$, a time integral operator $\int$ that is numerically evaluated by an ODE solver, and $\beta$ that calculates the time-averaged statistics from a time series of system states, i.e.,
$\begin{gathered}G(\theta):=\beta\Bigl{(}\bigl{\{}\bm{u}_{t_{0}}+\int_{t_{0}}^{t_{i}}\tilde{\mathcal{G}}(\tilde{\bm{u}}(\textbf{x},t),t;\theta)\mathop{}\!\mathrm{d}t\bigr{\}}_{i=0}^{l}\Bigr{)}.\end{gathered}$ (2.16)

### v3

@@ 2.2 Neural Operator
To construct the data-driven dynamical operator $\tilde{\mathcal{G}}$ in the continuous setting, i.e., viewing the system state $\bm{u}$ as a continuous spatial-temporal field instead of a discrete finite-dimensional vector, we rely on the recent developments of neural operators [28, 33]. As a high-level framework for modeling continuous dynamical systems, the neural dynamical operator can utilize any neural operator architecture to construct the data-driven dynamical operator $\tilde{\mathcal{G}}$. In this work, we present results using Fourier neural operator (FNO)  [28]. In general, neural operators can approximate $\mathcal{G}$ in Eq. (2.1) which is a non-linear mapping between infinite-dimensional spaces:
$\mathcal{G}:\mathcal{U}_{t}\times[0,T]\mapsto\mathcal{U}_{t},$ (2.5)
with a neural network $\tilde{\mathcal{G}}(\cdot;\bm{\theta})$ parameterized by $\bm{\theta}$. The key advantages of continuous mapping in Eq. (2.5) are: (i) the performance of the trained mapping is resolution-invariant, and (ii) the flexibility of using data with different discretizations. The resolution-invariant feature allows the model to learn from and evaluate the functions discretized in an arbitrary way. In other words, the model trained at one specific resolution can be used to perform inference at different resolutions of input and output functions, without a significantly increased error level. With the resolution-invariant property in the construction of the FNO architecture, the training process can also potentially combine the use of data in different resolutions.
We then focus on the use of FNO to construct the data-driven operator $\tilde{\mathcal{G}}$. The FNO applies a Fourier transform to convert the input spatial function from the spatial domain to the Fourier domain, processes the data within the Fourier domain, and subsequently performs an inverse Fourier transform to transform the results back to the spatial domain. Similarly, when the input function is defined on a temporal domain, it can still be transformed between the temporal domain and the Fourier domain. More specifically, FNO first linearly transforms the state function $\bm{u}(\cdot,t)\subset\mathcal{U}_{t}$ and temporal variable $t$ to lift the dimension from $\mathbb{R}^{d_{\bm{u}}+1}$ to $\mathbb{R}^{d_{\bm{v}}}$: $\bm{v}_{0}(\bm{x})=P(\bm{u}(\bm{x},t),t)$. Then $\bm{v}_{0}(\bm{x})$ will be iteratively transformed by Fourier layers with the output dimension staying the same as $d_{\bm{v}}$:
$\bm{v}_{n+1}(\bm{x})=\sigma(\bm{W}\bm{v}_{n}(\bm{x})+(\bm{K}\bm{v}_{n})(\bm{x})),\quad n=0,1\ldots N-1,$ (2.6)
@@
where $(\bm{K}\bm{v}_{n})(\bm{x})=\mathcal{F}^{-1}(\bm{R}\cdot(\mathcal{F}\bm{v}_{n}))(\bm{x})$ is the Fourier integral operator, $\mathcal{F}$ is Fourier transform and $\mathcal{F}^{-1}$ is its inverse, $\bm{R}$ is the complex-valued parameters in the neural network, $\bm{W}$ is a linear transformation, and $\sigma$ is a nonlinear activation function. At last, $\bm{v}_{N}(\bm{x})$ will be linearly transformed again to ensure the dimension of the final output same as the dynamics function in the space $\mathcal{U}_{t}$, i.e., $Q(\bm{v}_{N}(\bm{x}))=\frac{\partial\tilde{\bm{u}}(\bm{x},t)}{\partial t}\in\mathbb{R}^{d_{\bm{u}}}$.
@@
The classical loss function for training FNO quantifies the differences between the true dynamics $\frac{\partial\bm{u}(\bm{x},t)}{\partial t}$ and the predicted dynamics $\frac{\partial\tilde{\bm{u}}(\bm{x},t)}{\partial t}$. However, the data of $\frac{\partial\bm{u}(\bm{x},t)}{\partial t}$ may not be available in many applications, e.g., when the temporal resolution in the data of $\bm{u}(\bm{x},t)$ is not fine enough to obtain an informative estimation of its time derivative. Therefore, we may not be able to directly train the neural dynamical operator $\tilde{\mathcal{G}}$ by solving the optimization problem as typically done in the classical usage of FNO. To address this issue, we explore the training of a neural dynamical operator via the framework of neural ODE [45].
@@ 2.3 Neural ODE
The goal of neural ODE is to train the surrogate model $\tilde{\mathcal{G}}$, which is constructed via FNO in this work, to approximate $\mathcal{G}$ in Eq. (2.1) based on time series data $\{\bm{u}(t_{n})\}_{n=0}^{N}$ with $\bm{u}(t):=\bm{u}(\cdot,t)|D_{\bm{x}}^{(M)}$ which is defined in Section 2.1. With a given $\tilde{\mathcal{G}}$ and initial state of true system $\bm{u}(t_{0})$, the predictive system state at future time $t_{n}$ can be written as $\tilde{\bm{u}}(t_{n})=\bm{u}(t_{0})+\int_{t_{0}}^{t_{n}}\tilde{\mathcal{G}}(\tilde{\bm{u}}(t),t;\bm{\theta})\mathop{}\!\mathrm{d}t$, which can be obtained by an ODE solver in real applications. The trainable parameters $\bm{\theta}$ in $\tilde{\mathcal{G}}$ can then be obtained by minimizing the loss function in Eq. (2.3). With the loss function $J_{s}$ at each time step chosen as standard $\ell^{2}$-norm in this work, the total short-term loss for trajectory matching becomes
$L_{s}=\sum_{n=0}^{N_{s}}\|\bm{u}(t_{n})-\tilde{\bm{u}}(t_{n})\|^{2}.$ (2.7)
@@
The neural ODE framework [45] discussed two methods to calculate $\mathop{}\!\mathrm{d}L_{s}/\mathop{}\!\mathrm{d}\bm{\theta}$, including (i) backpropagation and (ii) adjoint sensitivity method, with which we can determine the optimal $\bm{\theta}$ by minimizing the short-term loss in Eq. (2.7) via gradient descent.
@@
Backpropagation is a classical gradient descent optimization method for training a neural network. One drawback of the backpropagation for neural ODE training is the memory cost to store the results of its forward pass, which is linearly proportional to the number of $\tilde{\mathcal{G}}$ evaluations. Although various types of memory management techniques [55] have been developed for automatic differentiation, the memory cost for matching a long time series data with a large scale model of $\tilde{\mathcal{G}}$ can still be inefficient or even infeasible.
@@
On the other hand, the adjoint sensitivity method calculates the gradient $\mathop{}\!\mathrm{d}L_{s}/\mathop{}\!\mathrm{d}\bm{\theta}$ by a forward and a backward ODE integration. The forward integration solves for the state $\tilde{\bm{u}}$ at future time step $t_{1},t_{2},...,t_{N}$, with the initial state $\tilde{\bm{u}}(t_{0})=\bm{u}(t_{0})$. By introducing an adjoint state $\bm{a}(t_{n})=\frac{\partial L_{s}}{\partial\tilde{\bm{u}}(t_{n})}$, the backward integration solves for an augmented states $[\tilde{\bm{u}}(t),\bm{a}(t),\frac{\partial L_{s}}{\partial\bm{\theta}}]^{\top}$ from $t_{n}$ to $t_{n-1}$, for $n=N,N-1,...,1$, with the initial state $\tilde{\bm{u}}(t_{N})$ evaluated from the forward integration, $\bm{a}(t_{N})=\frac{\partial L_{s}}{\partial\tilde{\bm{u}}(t_{N})}$, and $\frac{\partial L_{s}}{\partial\bm{\theta}}(t_{N})=\bm{0}$:
$\frac{\mathop{}\!\mathrm{d}\tilde{\bm{u}}(t)}{\mathop{}\!\mathrm{d}t}=\tilde{\mathcal{G}}(\tilde{\bm{u}}(t),t;\bm{\theta}),$ (2.8)
$\frac{\mathop{}\!\mathrm{d}\bm{a}(t)}{\mathop{}\!\mathrm{d}t}=-\bm{a}(t)^{\top}\frac{\partial\tilde{\mathcal{G}}(\tilde{\bm{u}}(t),t;\bm{\theta})}{\partial\tilde{\bm{u}}(t)},$
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\frac{\partial L_{s}}{\partial\theta}=-\bm{a}(t)^{\top}\frac{\partial\tilde{\mathcal{G}}(\tilde{\bm{u}}(t),t;\bm{\theta})}{\partial\bm{\theta}}.$
@@
It should be noted that the solved adjoint state $\bm{a}(t_{n})$ needs to be adjusted at $n=N-1,N-2,...,0$ during the backward integration, by adding a term of $\frac{\partial L_{s}}{\partial\tilde{\bm{u}}(t_{n})}$ to account for the fact that the loss function $L_{s}$ explicitly depends on the system state $\tilde{\bm{u}}(t_{n})$. During the whole backward integration, the vector-Jacobian products $\bm{a}^{\top}\frac{\partial\tilde{\mathcal{G}}}{\partial\tilde{\bm{u}}}$ and $\bm{a}^{\top}\frac{\partial\tilde{\mathcal{G}}}{\partial\bm{\theta}}$ can be evaluated on the fly without storing them in the memory. Therefore, the adjoint sensitivity method has a constant memory usage $O(1)$ with respect to the integration time steps, i.e., the number of $\tilde{\mathcal{G}}$ evaluations.
@@
For a short-term integration, the backpropagation often has no memory cost issue and is computationally faster than the adjoint method, considering that the Jacobian is stored and does not need to be evaluated on the fly in the backward pass. However, the memory cost issue would prevent the use of backpropagation if the data involves long-term integration, e.g., time-averaged statistics. In addition, gradient-based optimization can potentially encounter numerical issues (e.g., gradient blow-up) for the long-term information of chaotic systems. Therefore, backpropagation is used in this paper for short-term trajectory matching, while the long-term statistics of a chaotic dynamical system are incorporated by a derivative-free Kalman method, instead of the adjoint sensitivity method. More details about the derivative-free Kalman method can be found in Section 2.4.
@@ 2.4 Ensemble Kalman Inversion
The neural dynamical operator $\tilde{\mathcal{G}}$ in Eq. (2.2) can be calibrated to capture the long-term statistics of the true dynamical system by solving the optimization problem in Eq. (2.4) with ensemble Kalman inversion (EKI). Originated from ensemble Kalman filter [70, 71, 72], EKI [57, 73] has been developed as an optimization method to solve inverse problems. More specifically, the goal of EKI is to estimate the unknown parameters from noisy data in the general form of an inverse problem:
$\bm{y}=G(\bm{\theta})+\bm{\eta},$ (2.9)
where $\bm{y}$ is a vector of observation data, $G$ denotes a forward map, $\bm{\theta}$ corresponds to the trainable parameters, $\bm{\eta}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{\bm{\eta}})$ is often chosen as Gaussian random noises with a covariance matrix $\bm{\Sigma}_{\bm{\eta}}$. In this work, $\bm{y}=\beta(\{\bm{u}(t_{n})\}_{n=0}^{N_{l}})\in\mathbb{R}^{d_{\beta}}$ is the observed long-term statistics of the dynamical system, and the covariance matrix $\bm{\Sigma_{\bm{\eta}}}\in\mathbb{R}^{d_{\beta}\times d_{\beta}}$ is empirically chosen. In practice, the covariance matrix can also be determined by observing an ensemble of long-term statistics from the true system, with each ensemble member having a different initial condition. The forward map $G$ is a composition of data-driven dynamical operator $\tilde{\mathcal{G}}$, a time integral operator $\int$ that is numerically evaluated by an ODE solver, and $\beta$ that calculates the time-averaged statistics from a time series of system states, i.e.,
$\begin{gathered}G(\bm{\theta}):=\beta(\{\tilde{\bm{u}}(t_{n})\}_{n=0}^{N_{l}})=\beta\Bigl{(}\bigl{\{}\bm{u}(t_{0})+\int_{t_{0}}^{t_{n}}\tilde{\mathcal{G}}(\tilde{\bm{u}}(\bm{x},t),t;\bm{\theta})\mathop{}\!\mathrm{d}t\bigr{\}}_{n=0}^{N_{l}}\Bigr{)}.\end{gathered}$ (2.10)
To match the observed data $\bm{y}$ and the output of forward map $G$, EKI is employed to identify the optimal parameters $\bm{\theta}$. The optimization problem in Eq. (2.4) solved by EKI can be written as
$\min_{\bm{\theta}}||\bm{\Sigma}_{\bm{\eta}}^{-\frac{1}{2}}(\bm{y}-G(\bm{\theta}))||^{2}.$ (2.11)
To solve the optimization problem in Eq. (2.11), the EKI formula that iteratively updates the ensemble estimation of parameters $\{\bm{\theta}^{(j)}\}_{j=1}^{J}$ is:
$\begin{gathered}\bm{\theta}_{n+1}^{(j)}=\bm{\theta}_{n}^{(j)}+\bm{\Sigma}_{n}^{\bm{\theta}\bm{g}}(\bm{\Sigma}_{n}^{\bm{gg}}+\bm{\Sigma}_{\bm{\eta}})^{-1}(\bm{y}^{(j)}-\bm{g}_{n}^{(j)}),\end{gathered}$ (2.12)
where the index $n$ denotes the $n$-th EKI step and $\bm{y}^{(j)}$ corresponds to the perturbed observation data $\bm{y}$ via sampling the noises $\bm{\eta}$. With the ensemble of parameters $\{\bm{\theta}_{n}^{(j)}\}_{j=1}^{J}$ at the $n$-th EKI step, the terms $\bm{g}_{n}^{(j)}$, $\bm{\Sigma}_{n}^{\bm{\theta g}}$, and $\bm{\Sigma}_{n}^{\bm{gg}}$ in Eq. (2.12) are calculated as:
$\begin{gathered}\bar{\bm{\theta}}_{n}=\frac{1}{J}\sum_{j=1}^{J}\bm{\theta}_{n}^{(j)},~{}~{}\bm{g}^{(j)}_{n}=G(\bm{\theta}_{n}^{(j)}),~{}~{}\bar{\bm{g}}_{n}=\frac{1}{J}\sum_{j=1}^{J}\bm{g}_{n}^{(j)},\\ \bm{\Sigma}_{n}^{\bm{\theta}\bm{g}}=\frac{1}{J-1}\sum_{j=1}^{J}(\bm{\theta}_{n}^{(j)}-\bar{\bm{\theta}}_{n})(\bm{g}_{n}^{(j)}-\bar{\bm{g}}_{n})^{\top},\\ \bm{\Sigma}_{n}^{\bm{gg}}=\frac{1}{J-1}\sum_{j=1}^{J}(\bm{g}_{n}^{(j)}-\bar{\bm{g}}_{n})(\bm{g}_{n}^{(j)}-\bar{\bm{g}}_{n})^{\top}.\end{gathered}$ (2.13)